show that unit vector normal to the surface n= ∇φ ÷ | ∇φ | . find n for the ellipsoid.


φ = a(x^2) + b(y^2) +c(z^2)

1
Expert's answer
2022-01-23T16:14:57-0500

Given:

"\u03c6 = ax^2 + by^2+cz^2"


The gradient of scalar field is given by

"\nabla \varphi=\frac{\partial \varphi}{\partial x}\hat i+\frac{\partial \varphi}{\partial y}\hat j+\frac{\partial \varphi}{\partial z}\hat k"

Hence

"\nabla \varphi=2ax\hat i+2by\hat j+2cz\hat k"

"|\nabla \varphi|=2\sqrt{(ax)^2 +(by)^2+(cz)^2}"

"{\vec n}=\frac{\nabla \varphi}{|\nabla \varphi|}=\frac{ax\hat i+by\hat j+cz\hat k}{\sqrt{(ax)^2 +(by)^2+(cz)^2}}"